OpenStudy (anonymous):

Which number is IRRATIONAL? A) Square root of 25 B) Sqauare root of 32 C) Square root of 49 D) Square root of 64

OpenStudy (anonymous):

\[\sqrt{25} = 5\] \[\sqrt{32} = 5.656854...\] \[\sqrt{49} = 7\] \[\sqrt{64} = 8\] Irrational number is a number that cannot be expressed as an integer or a fraction. (an integer is any number -1, 0, 1, 2 etc. The correct answer is B) Sqauare root of 32